OpenStudy (anonymous):

A neighborhood has provided trash cans to residents to dispose of their waste. All trash cans have a diameter of 0.5 meters and a height of 1.5 meters. What is the minimum number of trash cans required to dispose of 10 cubic meters of waste?

OpenStudy (anonymous):

A. 34 B. 101 C. 33 D. 102

OpenStudy (cwrw238):

first calculate the volume of each trash can can using v = pi r^2 h where r = 1/2 * 0.5 and h = 1.5 then divide this into 10 cu meters - then round up to nearest whole number

OpenStudy (cwrw238):

any questions?

OpenStudy (anonymous):

I got it thank you!

OpenStudy (cwrw238):

yw